Ms. Imo, who is single, purchased her first home in 1991 for $85,000 and sold it in May 2000 for $178,500. She purchased her second home in July 2000 for $385,000 and sold it this year for $700,000.

A. Compute Ms. Imo’s taxable gain on the 2000 sale and on this year’s sale.

B. Compute the income tax and Medicare contribution tax on her gain this year if her preferential rate on long-term capital gain is 20 percent.
Answer #1

a)

Ms. Imo realized a gain on 2000 sale and excluded the entire gain from income

= 178,500-85,000

= 93,500

Ms. imo realized gain amount on the current year sale

= 700,000-385,000

= 315,000

Ms. Imo excluded 250,000 gain from income and recognized taxable gain

= 65,000

b)

Ms. Imo's 65,000 taxable gain is long-term capital gain, and her tax

= 65,000*20%

= 13,000